Given that A= 3 ‐1 2 and B= 4 [‐1 2]

[5 1 7] 5 1 3

I. Evaluate 3A and 2B

ii. 3A‐2B

Given that C= 1 6

[3 9]

4 -3 Evaluate CB



1
Expert's answer
2021-12-07T11:28:12-0500
"A=\begin{bmatrix}
 3 & -1 & 2\\
 5 & 1& 3
\end{bmatrix}, B=\begin{bmatrix}
 4 & -1 & 2\\
 5 & 1 & 7
\end{bmatrix}"

"C=\begin{bmatrix}
 1 & 6\\
 3 & 9 \\
 4 & -3
\end{bmatrix}"

i.


"3A=\begin{bmatrix}
 9 & -3 & 6\\
 15 & 3 & 9
\end{bmatrix}"

"2B=\begin{bmatrix}
 8 & -2 & 4\\
 10 & 2 & 14
\end{bmatrix}"


ii.


"3A-2B=\begin{bmatrix}
 1 & -1 & 2\\
 5 & 1 & -5
\end{bmatrix}"

iii.


"CB=\begin{bmatrix}
 1 & 6\\
 3 & 9 \\
 4 & -3
\end{bmatrix}\begin{bmatrix}
 4 & -1 & 2\\
 5 & 1 & 7
\end{bmatrix}"

"=\begin{bmatrix}
 1(4)+6(5) & 1(-1)+6(1) & 1(2)+6(7)\\
 3(4)+9(5) & 3(-1)+9(1) & 3(2)+9(7) \\
 4(4)-3(5) & 4(-1)-3(1) & 4(2)-3(7)
\end{bmatrix}"

"=\begin{bmatrix}
 34 & 5 & 44\\
 57 & 6 & 69\\
 1 & -7 & -13\\
\end{bmatrix}"